Please, ia m not even sure if i am in the right area. Wont someone help me please?????
Solve the system by the elimination method of these:
1) 5x + 5y = -11
7x - 2y = 13
2) 0.3x - 0.2y = 4
0.2x + 0.5y = 1
3) 6x - 9y = 13.5
7y - 2x = -6.5
4) 0.05x + 0.25y = 11
0.15x + 0.05y = 12
On these it says "What is the solution of the system". (Type an order pair. Type an integer or a decimal. Type N if there is no solution. Type I if there are infinitely many solutions)
Will someone, anyone, please help me. I am so confused. thank you in advance to anyone that will take the time to help.

Answer by stanbon(75887)

Solve the system by the elimination method of these:
1)
5x + 5y = -11
7x - 2y = 13
-------------------
Multiply thru the 1st equation by 7; Multiply thru the 2nd equation by 5 to get:
35x + 35y = -77
35x - 10y = 65
-----------------
Subtract the 2nd equation from the 1st so you can solve for "y":
25y = -142
y = -142/25
--------------------
Substitute that into 5x+5y=-11 so you can solve for "x":
5x + 5(-142/25) = -77
5x - (142/5) = -77
5x = -77 + (142/5)
5x = -243/5
x = -243/25
=======================
2) 0.3x - 0.2y = 4
0.2x + 0.5y = 1
Multiply thru both equations by 10 to get rid of the decimals:
---------------------
3x - 2y = 40
2x + 5y = 10
-----------------
Multiply the 1st equation by 2; Multiply the 2nd equation by 3 to get:
6x -4y = 80
6x +15y = 30
------------------
Subtract the 1st equation from the 2nd to get:
19y = -50
y = -50/19
------------------
Substitute that into 3x-2y=40 to solve for "x":
3x-2(-50/19) = 40
3x + (100/19) = 40
3x = 660/19
x = 220/19
=======================
Comment: Maybe you can handle #'s 3 and 4 now.
Cheers,
Stan H.


3)
6x - 9y = 13.5
7y - 2x = -6.5
Ans: x = -57/34; y=-89/34
==========================

4)
0.05x + 0.25y = 11
0.15x + 0.05y = 12
Ans: x=70 ; y = 30
==========================
Cheers,
Stan H.